# Bayesian Multifractal Image Segmentation

### v1

@@ 5.4 Gibbs Sampler and Bayesian Estimators
for $j=j_{1},...,j_{2}$. Note that the computed labels are downsampled by a factor $2^{j_{1}}$ compared to the input image. In practice, the labels computed at the finest scale $\hat{\mathbf{z}}_{j_{1}}\in\mathbb{R}^{N_{1}\times N_{1}}$ are interpolated (using, e.g, nearest neighbors) up to $N\times N$ to obtain labels for each pixel of the image, see [22] for alternative strategies. Finally, the updates for the granularity parameter vector $\boldsymbol{\beta}$ are performed only during the burn-in period of the sampler, as suggested in [40], in order to reduce the computational complexity. After the burn-in period, $\boldsymbol{\beta}$ is fixed to its current value.
@@
[…]
$K:$ number of regions
@@
       Compute Fourier coefficients x=[$\textbf{x}=[\textbf{x}^{T}_{1},...,\textbf{x}^{T}_{K}]^{T}$ using $\textbf{z}^{(t-1)}$ via (16). Sample $\textbf{z}^{(t)}\sim p\Big{[}\textbf{z}|\boldsymbol{\Theta}^{(t-1)},\textbf{z}^{(t-1)},\boldsymbol{\beta}^{(t-1)},\textbf{x},\boldsymbol{\mu}\Big{]}$ via (29). Sample $\boldsymbol{\Theta}^{(t)}\sim f\big{(}\boldsymbol{\Theta}|\boldsymbol{\Theta}^{(t-1)},\textbf{z}^{(t)},\boldsymbol{\beta}^{(t-1)},{\textbf{x}},\boldsymbol{\mu}\big{)}$ and $\boldsymbol{\mu}$ via (31). if $t<N_{b}$ then
@@ 5.6 Computational complexity
The computational complexity of the proposed approach is dominated by that of the Fourier transforms. Since $N_{j}\propto N$ for all $j$, the complexity of the multifractal parameter computation requires $\mathcal{O}(N^{2}\log N)$ operations, performed $K$ times. The cost of constructing the debiasing matrix in (19) is $\mathcal{O}(N^{2})$. The multiscale Potts model update also requires $\mathcal{O}\big{(}N^{2}\big{)}$ operations to compute the probability of wavelet leaders to belong to each class. Consequently, the final complexity of the proposed method is $\mathcal{O}\big{(}KN^{2}\log N\big{)}$.
@@ 6 Numerical Experiments
The proposed joint Bayesian segmentation and parameter estimation approach is first evaluated on synthetic 2D images constructed using multifractal random walks (MRWs) by means of Monte Carlo simulations. Some results obtained using real-world data are then presented. MRWs are non-Gaussian processes with stationary increments and are constructed in order to have multifractal properties mimicking those of Mandelbrot’s log-normal cascades, see [42] for details.
@@ 6.1.1 Synthetic Data
To validate the proposed method under controlled conditions, two scenarios with known ground-truth labels and multifractal parameters were considered. Specifically, the first scenario has $K=2$ different multifractal regions: a background with multifractality parameter $\theta^{1}_{1}=-0.02$ and a centered disk with parameter $\theta^{2}_{1}=-0.08$, whereas the second scenario has $K=3$ regions: a background with $\theta^{1}_{1}=-0.02$ and two disks with parameters $\theta^{2}_{1}=-0.08$ and $\theta^{3}_{1}=-0.16$. The experiments were run on 100 independent realizations of MRW with these parameters. An illustration of a single realization of an image for each scenario and their corresponding true masks is displayed in Figs. 3 (a) and (b), respectively.
@@ 6.1.3 Parameter Settings
The Daubechies mother wavelet with $N_{\psi}=1$ vanishing moments is used in the 2D DWT. The hyperparameters of the independent priors for the multifractal parameters were set to $(\alpha_{i,k},\gamma_{i,k})=(10^{-3},10^{-3})$ [4]. The set of scales used for the experiment with $K=2$ classes was set to $(j_{1},j_{2})=(1,3)$, whereas $(j_{1},j_{2})=(1,2)$ for $K=3$. The number of iterations for all the experiments was set to $N_{m}=300$, with a burn-in period of $N_{b}=30$.
@@ 6.2 Estimation Performance
6.2.1 Multifractal Parameter Estimation on an Irregular Lattice
The proposed approach is first evaluated in a setting where the labels are known and fixed, specifically for the case where $K=2$, in order to assess the estimation performance of the multifractal parameter on an irregular grid. Furthermore, to illustrate the impact of the multifractality parameter value on the image segmentation, the multifractality value in the disk of the MRWs was varied in a range of weak (close to zero) to strong values $\theta^{2}_{1}\in\{-0.005,-0.08,-0.2\}$, fixing the background parameter to $\theta^{1}_{1}=-0.02$. Results are reported in Table 1. They indicate that the proposed approach yields satisfactory performance (small RMSE values), comparable to that obtained for regular lattices [4, 29].
@@ 6.2.2 Multifractal Parameter and Label Estimation
Table 2 shows the performance of the multifractal parameter estimation when the labels and granularity parameters are also unknown and estimated jointly, following the previously described setting. The results show that the RMSE values slightly increase, due to the larger number of unknowns to be estimated, but that overall performance remains comparable to Table 1. Table 3 reports the DSC values and the segmentation errors for the two experiments, allowing the label estimation performance to be evaluated. The results indicate that the segmentation error decreases as the difference in the multifractal parameter between the regions increases. This behavior is indeed expected, since the increased parameter contrast should facilitate segmentation. Overall, the segmentation performance is very satisfactory, except for the extreme case $\theta^{2}_{1}=-0.005$ that corresponds to nearly no multifractality.
@@ 6.3 Comparison to the State of the Art
The proposed segmentation approach is compared to different unsupervised state-of-the-art methods for texture///image segmentation, including deep learning methods, classical texture segmentation, and the classical $k$-means clustering method, a benchmark in segmentation, used to estimate labels for different features, including multifractal, pixel intensity, and multiscale features. Specifically:
  • The $k$-means on multifractal estimates (Kmeans-MF) from regression: the initialization of the proposed method described in 5.5.
  • The $k$-means on a localized Bayesian multifractal estimates (Kmeans-GaMRF) [4]: makes use of a smoothing gamma Markov random field (GaMRF) prior [^2] Toolbox available at https://www.irit.fr/~Herwig.Wendt/CIMPA2017.html for the multifractal estimates as features for multifractal-based segmentation.
@@
  • The $k$-means on Gabor filter features (Kmeans-Gabor) [11]: a benchmark for classical multiscale features extraction without multifractal model.
  • The factorization-based texture segmentation (FSeg) method [16]: computes multiscale Gabor coefficients followed by a feature selection procedure by a matrix factorization step. This method is selected given its multiscale nature for segmentation. The hyperparameters are configured as in [16].
  • The segment anything (SAM) method [12]: it relies on a promptable segmentation that takes as inputs the image and a handcrafted reference (such as points, boxes or mask logits), and returns a segmentation mask [12]. The pre-trained SAM model was supervised by providing spatial coordinates of two pixels randomly selected from each region.
  • The Mask R-CNN architecture [13]: a pre-trained convolutional neural network for instance segmentation that generates a pixel-level binary mask for each detected object in the given image. A threshold of 0.98 was applied to the network’s output masks for the detected objects.
The SAM and Mask R-CNN models were selected since they are outstanding methods of the state of the art in deep-based image/instance segmentation.
Figure 3 displays the segmentation results obtained using the different methods for a single MRW realization, when $K=2$ (top row) and $K=3$ (bottom row). The proposed method achieves an average DSC of $\approx 0.85$, significantly outperforming the other methods, and provides visually satisfactory results. All other baselines yield significantly worse results, illustrating the difficulty of segmenting multifractal images. Tables 4 and 5 provide a more quantitative and systematic comparison in terms of average performance for 100 independent realizations of MRW for $K=2$ and $3$, respectively. It can be seen that the proposed method significantly outperforms the other methods, showing overall best segmentation scores in terms of DSC and Error by a large margin.
Table 4: Segmentation baseline comparison for the label estimation when ${K=2}$ in terms of average DSC and segmentation Error percentage of 100 realizations, and its the standard deviation is shown in parenthesis. The best scores are highlighted in bold.
@@
Kmeans-GaMRF $0.843\,({0.027})$ $0.567\,(0.097)$ $22.8\,(3.7)$
Kmeans-MF $0.745\,(0.079)$ $0.539\,(0.111)$ $32.2\,(7.9)$
Kmeans-Feat $0.643\,(0.071)$ $0.383\,(0.088)$ $44.2\,(5.7)$
Kmeans-GABOR $0.683\,(0.084)$ $0.443\,(0.126)$ $39.6\,(8.4)$
FSeg $0.637\,(0.050)$ $0.416\,(0.077)$ $44.6\,(5.2)$
Mask-RCNN $0.813\,(0.042)$ $0.300\,(0.250)$ $28.5\,({0.1})$
SAM $0.196\,(0.215)$ $0.458\,({0.054})$ $62.7\,(8.8)$
Table 5: Average DSC for 100 realizations of the baseline comparison when ${K=3}$. The standard deviation is shown in parenthesis.
@@
Kmeans-GaMRF $\textbf{0.846}\,{(0.051)}$ $0.328\,(0.124)$ $0.439\,(0.173)$
Kmeans-MF ${0.669}\,(0.094)$ $0.230\,(0.083)$ $0.390\,(0.147)$
Kmeans-Feat ${0.505}\,(0.150)$ $0.219\,(0.068)$ $0.139\,(0.078)$
Kmeans-GABOR ${0.515}\,(0.091)$ $0.353\,(0.102)$ $0.210\,(0.119)$
FSeg ${0.557}\,(0.118)$ $0.322\,(0.113)$ $0.198\,(0.141)$
Mask-RCNN ${0.765}\,(0.292)$ $0.093\,(0.166)$ $0.037\,(0.112)$
SAM ${0.547}\,(0.186)$ $0.282\,(0.112)$ $0.016\,(0.088)$
6.4 Illustration for Real data: Tumor Segmentation
This section studies the effectiveness of the proposed approach for real data. Specifically, the proposed segmentation method is applied to a mammogram for tumor segmentation in breast screening. Mammography remains the primary technique for detecting and diagnosing breast cancer, and early detection can significantly reduce treatment costs. Therefore, the development of methodologies that enable accurate tumor detection is of great interest, with multifractal-based methods showing promising results [43]. In this context, the proposed method works under the assumption that tumor and normal tissue are characterized by different multifractal parameters, and $K=2$.
6.4.1 Dataset
The mammography image used for this experiment was taken from the Mini-DDSM (Digital Database of Screening Mammography) dataset [^3] Mammography dataset available at https://ardisdataset.github.io/MiniDDSM/. The set used for the simulation is “cancer/1727”.. In this dataset, each image has a segmentation label, or contour, that delineates the margins of the tumor [^4] It is important to note that the provided contour serves just as a reference to localize the tumor, as literature has reported that such contours may exhibit inaccuracies. Nevertheless, they offer a sufficiently accurate approximation for characterizing the tumor. For comparison purposes, a segmentation mask is computed from the provided contour by performing a flood-fill operation on the binary mask.. The mammography image is cropped to retain only the region of interest shown in Figure 4-(a), removing the surrounding background.
@@ 6.4.1 Dataset
Figure 4: Comparison of segmentation results on a mammography image for tumor detection, with the tumor contour shown in red on the resulting masks and the segmentation error displayed below each result. Notably, SAM method achieves the best segmentation performance, yet it requires ground-truth pixels from each region as prompts for its operation, making it supervised. The second best segmentation is achieved with the proposed method, which is fully unsupervised, showing its capability of modeling different multifractilities on an image for segmentation purposes.
6.4.2 Detection Performance

### v2

@@ 5.4 Gibbs Sampler and Bayesian Estimators
for $j=j_{1},...,j_{2}$. Note that the computed labels are downsampled by a factor $2^{j_{1}}$ compared to the input image. In practice, the labels computed at the finest scale $\hat{\mathbf{z}}_{j_{1}}\in\mathbb{R}^{N_{1}\times N_{1}}$ are interpolated (using, e.g, nearest neighbors) up to $N\times N$ to obtain labels for each pixel of the image, see [choi2001multiscale] for alternative strategies. Finally, the updates for the granularity parameter vector $\boldsymbol{\beta}$ are performed only during the burn-in period of the sampler, as suggested in [pereyra2014maximum], in order to reduce the computational complexity. After the burn-in period, $\boldsymbol{\beta}$ is fixed to its current value.
@@
Input: $\textbf{F},K:$ input image, number of regions
@@
  Compute Fourier coefficients $\textbf{x}=[\textbf{x}^{T}_{1},...,\textbf{x}^{T}_{K}]^{T}$ using $\textbf{z}^{(t-1)}$ via (16). Sample $\textbf{z}^{(t)}\sim p\Big[\textbf{z}|\boldsymbol{\Theta}^{(t-1)},\textbf{z}^{(t-1)},\boldsymbol{\beta}^{(t-1)},\textbf{x},\boldsymbol{\mu}\Big]$ via (29). Sample $\boldsymbol{\Theta}^{(t)}\!\sim\!f\big(\boldsymbol{\Theta}|\boldsymbol{\Theta}^{(t-1)},\!\textbf{z}^{(t)},\boldsymbol{\beta}^{(t-1)}\!,{\textbf{x}},\boldsymbol{\mu}\big)$, and $\boldsymbol{\mu}$ via (31). if $t<N_{b}$ then
@@ 5.6 Computational complexity
The computational complexity of the proposed approach is dominated by that of the Fourier transforms. Since $N_{j}\propto N$ for all $j$ (where $\propto$ means “proportional to”), the estimation of the multifractal parameter requires $\mathcal{O}(N^{2}\log N)$ operations, performed $K$ times. The cost of constructing the debiasing matrix in (19) is $\mathcal{O}(N^{2})$. The multiscale Potts model update also requires $\mathcal{O}\big(N^{2}\big)$ operations to compute the probability of wavelet leaders to belong to each class. Consequently, the final complexity of the proposed method is $\mathcal{O}\big(KN^{2}\log N\big)$. Although demanding for high-resolution images, computational efficiency can be improved by implementing the debiasing matrix at a lower level (e.g., C/C++) and parallelizing the multiscale Potts updates on a GPU.
@@ 6 Numerical Experiments
The proposed joint Bayesian segmentation and parameter estimation approach is first evaluated on synthetic 2D images constructed using multifractal random walks (MRWs) by means of Monte Carlo simulations [^2] The source code is available at https://github.com/kareth-leon/MFA_Seg. Some results obtained using real-world data are then presented. MRWs are non-Gaussian processes with stationary increments and are constructed in order to have multifractal properties mimicking those of Mandelbrot’s log-normal cascades, see [robert2010gaussian] for details.
@@ 6.1.1 Synthetic Data
To validate the proposed method under controlled conditions, two scenarios with known ground-truth labels and multifractal parameters were considered. Specifically, the first scenario has $K=2$ different multifractal regions: a background with multifractality parameter $\theta^{1}_{1}=-0.02$ and a centered disk with parameter $\theta^{2}_{1}=-0.08$, whereas the second scenario has $K=3$ regions: a background with $\theta^{1}_{1}=-0.02$ and two disks with parameters $\theta^{2}_{1}=-0.08$ and $\theta^{3}_{1}=-0.16$. The experiments were run on 100 independent realizations of MRWs with these parameters, where the computation time for one realization is in average 58 seconds [^3] The simulations were conducted in a desktop architecture with an Intel Corei7, 16 GB RAM, using MATLAB R2024b.. An illustration of a single realization of an image for each scenario and their corresponding true masks is displayed in Figs. 4 (a) and (b), respectively.
@@ 6.1.3 Parameter Settings
The Daubechies mother wavelet with $N_{\psi}=1$ vanishing moments is used in the 2D DWT. The hyperparameters of the independent priors for the multifractal parameters were set to $(\alpha_{i,k},\gamma_{i,k})=(10^{-3},10^{-3})$ [wendt2018multifractal]. In practice, these constants are fixed to small values, ensuring that they are close to non-informative Jeffreys priors [wendt2018multifractal, pereyra2013estimating]. The set of scales used for the experiment with $K=2$ classes was set to $(j_{1},j_{2})=(1,3)$, whereas $(j_{1},j_{2})=(1,2)$ for $K=3$. Note that $j_{2}$ is limited by image dimensions ($j_{2}<\log_{2}N$), nonetheless, in this study, $j_{2}$ was set between 2 and 3, as the coarsest scales were found to retain meaningful pixel-level information. The number of iterations for all the experiments was set to $N_{m}=300$, with a burn-in period of $N_{b}=30$, since it showed to be sufficient for sampling convergence (see next subsection for convergence details).
@@ 6.2 Estimation Performance
6.2.1 Multifractal Parameter Estimation on Irregular Shapes
The proposed approach is first evaluated in a setting where the labels are known and fixed with $K=2$ classes, in order to assess the estimation performance of the multifractal parameter on an irregular domain. Furthermore, to illustrate the impact of the multifractality parameter value on the image segmentation, the multifractality value in the disk of the MRWs was varied in a range of weak (close to zero) to strong values $\theta^{2}_{1}\in\{-0.005,-0.08,-0.2\}$, fixing the background parameter to $\theta^{1}_{1}=-0.02$. Results are reported in Table 1. They indicate that the proposed approach yields satisfactory performance (small RMSE values), comparable to that obtained for regular shapes [wendt2018multifractal, combrexelle2015bayesian].
@@ 6.2.2 Multifractal Parameter and Label Estimation
Table 2 shows the performance of the multifractal parameter estimation when the labels and granularity parameters are also unknown and estimated jointly, following the previously described setting. The results show that the RMSE values slightly increase, due to the larger number of unknowns to be estimated, but that the overall performance remains comparable to Table 1. Table 3 reports the DSC values and the segmentation errors for the two experiments, allowing the label estimation performance to be evaluated. The results indicate that the segmentation error decreases as the difference in the multifractal parameter between the regions increases. This behavior is indeed expected, since the increased parameter contrast should facilitate segmentation. Overall, the segmentation performance is very satisfactory, except for the extreme case $\theta^{2}_{1}=-0.005$ that corresponds to nearly no multifractality.
6.2.3 Sampler Convergence
To evaluate the convergence of the proposed method, we use the potential scale reduction factor (PSRF), a convergence diagnostic for Monte Carlo simulations [gelman1992inference]. It compares the variance within the chain to the variance between chains in multiple chains, where PSRF <1.2<1.2 indicates good convergence of the sampler (see [gelman1992inference, dobigeon2007joint] for further details). Any parameter sampled in the Monte Carlo simulations can serve as a reference to monitor convergence. For illustration purposes, we chose to monitor the multifractal value of the background of Fig. 4-a, $\theta^{1}_{1}=-0.02$, during 1000 iterations. Figure 5 shows five chains of the parameter with different initial values. It can be seen that the chains converge to similar values, showing that the method converges even when changing the initial value of the chain. The PSRF value obtained for this parameter with $1000$ Monte Carlo runs is $1.0009$, exhibiting the convergence of the sampler.
Figure 5: Convergence assessment with five Markov chains with different initializations.
@@ 6.3 Comparison to the State of the Art
The proposed segmentation approach is compared to unsupervised state-of-the-art methods for texture//image segmentation, including deep learning methods, classical texture segmentation, and the classical $k$-means clustering method, a benchmark in segmentation, used to estimate labels for different features, including multifractal, pixel intensity, and multiscale features. These methods are summarized below:
  • The $k$-means on multifractal parameters estimated by regression (Kmeans-MF). Note that this method was used to initialize the suggested method described in 5.5.
  • The $k$-means on a localized Bayesian multifractal estimates (Kmeans-GaMRF) [wendt2018multifractal], which makes use of a smoothing gamma Markov random field (GaMRF) prior [^4] Toolbox available at https://www.irit.fr/~Herwig.Wendt/CIMPA2017.html. The resulting multifractal estimates are used as features for $k$-means segmentation.
@@
  • The $k$-means on Gabor filter features (Kmeans-Gabor) [jain1991unsupervised]: a benchmark for classical multiscale features extraction without multifractal model.
  • The factorization-based texture segmentation (FSeg) method [yuan2015factorization]: computes multiscale Gabor coefficients followed by a feature selection procedure by a matrix factorization step. This method is selected given its multiscale nature for segmentation. The hyperparameters were adjusted as in [yuan2015factorization].
  • The segment anything (SAM) method [kirillov2023segment]: it relies on a promptable segmentation that takes as inputs the image and a handcrafted reference (such as points, boxes or mask logits), and returns a segmentation mask [kirillov2023segment]. The pre-trained SAM model was supervised by providing the spatial coordinates of two pixels randomly selected from each region.
  • The Mask R-CNN architecture [he2017mask]: a pre-trained convolutional neural network for instance segmentation that generates a pixel-level binary mask for each detected object in the given image. A threshold of 0.98 was applied to the network’s output masks for the detected objects.
The SAM and Mask R-CNN models were selected since they are outstanding methods of the state of the art in deep-based image/instance segmentation. Additionally, note that the SAM model relies on multiple transformers and attention modules that model long-range dependencies to handle global context and complex scenes [kirillov2023segment].
Figure 4 displays the segmentation results obtained using the different methods for a single MRW realization, when $K=2$ (top row) and $K=3$ (bottom row). The proposed method yields an average DSC close to $0.85$, significantly outperforming the other methods, and provides visually satisfactory results. All other baselines lead to significantly worse results, illustrating the difficulty of segmenting multifractal images. Tables 4 and 5 provide a more quantitative and systematic comparison in terms of average performance for 100 independent realizations of MRW for $K=2$ and $3$. It can be seen that the proposed method significantly outperforms the other methods, showing overall best segmentation scores in terms of DSC and Error by a large margin.
Table 4: Segmentation baseline comparison for the label estimation when K=2{K=2} in terms of average DSC and segmentation Error percentage of 100 realizations, the standard deviation is shown in parenthesis. The best scores are highlighted in bold.
@@
Kmeans-GaMRF 0.843(0.027)0.843\,({0.027}) 0.567(0.097)0.567\,(0.097) 22.8(3.7)22.8\,(3.7)
Kmeans-MF 0.745(0.079)0.745\,(0.079) 0.539(0.111)0.539\,(0.111) 32.2(7.9)32.2\,(7.9)
Kmeans-Feat 0.643(0.071)0.643\,(0.071) 0.383(0.088)0.383\,(0.088) 44.2(5.7)44.2\,(5.7)
Kmeans-GABOR 0.683(0.084)0.683\,(0.084) 0.443(0.126)0.443\,(0.126) 39.6(8.4)39.6\,(8.4)
FSeg 0.637(0.050)0.637\,(0.050) 0.416(0.077)0.416\,(0.077) 44.6(5.2)44.6\,(5.2)
Mask-RCNN 0.813(0.042)0.813\,(0.042) 0.300(0.250)0.300\,(0.250) 28.5(0.1)28.5\,({0.1})
SAM 0.196(0.215)0.196\,(0.215) 0.458(0.054)0.458\,({0.054}) 62.7(8.8)62.7\,(8.8)
Table 5: Average DSC for 100 realizations of the baseline comparison when K=3{K=3}. The standard deviation is shown in parenthesis.
@@
Kmeans-GaMRF $\textbf{0.846}\,{(0.051)}$ 0.328(0.124)0.328\,(0.124) 0.439(0.173)0.439\,(0.173)
Kmeans-MF 0.669(0.094){0.669}\,(0.094) 0.230(0.083)0.230\,(0.083) 0.390(0.147)0.390\,(0.147)
Kmeans-Feat 0.505(0.150){0.505}\,(0.150) 0.219(0.068)0.219\,(0.068) 0.139(0.078)0.139\,(0.078)
Kmeans-GABOR 0.515(0.091){0.515}\,(0.091) 0.353(0.102)0.353\,(0.102) 0.210(0.119)0.210\,(0.119)
FSeg 0.557(0.118){0.557}\,(0.118) 0.322(0.113)0.322\,(0.113) 0.198(0.141)0.198\,(0.141)
Mask-RCNN 0.765(0.292){0.765}\,(0.292) 0.093(0.166)0.093\,(0.166) 0.037(0.112)0.037\,(0.112)
SAM 0.547(0.186){0.547}\,(0.186) 0.282(0.112)0.282\,(0.112) 0.016(0.088)0.016\,(0.088)
6.4 Experiments using real data
This section studies the effectiveness of the proposed approach for real data segmentation. The proposed segmentation method is first applied to a mammogram image for tumor segmentation in breast screening , and second, to a synthetic aperture radar (SAR) image for water body detection.
Mammography remains the primary technique for detecting and diagnosing breast cancer. Early cancer detection can significantly reduce treatment costs. Therefore, the development of methodologies that enable accurate tumor detection is of great interest, with promising results obtained using multifractal-based methods [lopes2009fractal]. On the other hand, water body segmentation based on SAR images has become a research hotspot in remote sensing applications given that SAR satellites are able to operate any time, regardless of weather or lighting conditions [kim2021large, chen2023towards]. However, SAR images are contaminated by multiplicative speckle noise, making their use difficult for computer vision tasks [chen2023towards]. From these scenarios, it is important to remark that the applicability of the proposed method can be determined by assessing whether the data exhibit multifractal, scale-free behavior. This can be verified by checking if the scaling relation (3) [combrexelle2015bayesian] holds empirically, which is the case for the selected scenarios.
@@ 6.4 Experiments using real data
Figure 6: Comparison of segmentation results on a mammography image for tumor detection, with the tumor contour shown in red on the resulting masks and the segmentation error displayed below each result. Notably, SAM method achieves the best segmentation performance, yet it requires ground-truth pixels from each region as prompts for its operation, making it supervised. The second best segmentation is achieved with the proposed method, which is fully unsupervised, showing its capability of modeling different multifractalities on an image for segmentation purposes.
6.4.1 Tumor segmentation